the density of neon will be highest at


(a)STP


(b)0degree,2atm


(c)273degree,1atm


(d) 273degree,2atm

Grade:12

3 Answers

rajan jha
13 years ago

DENSITY(d)= pM/RT,  i.e, it is directly proportional yo pressure and inversely proportional to temperatutre. Thus the condition is highest pressure and low temperature

like my answer

vikas askiitian expert
13 years ago

PV =nRT            ( n is number of moles of gas)

n = mass of gas /M.wt   & dencity(d) is mass/V

PV=(mass) RT/M.wt

P=dRT/M.wt         or

d= p(M.wt)/RT

from above eq , if the pressure is more and temprature is less then dencity will be high.....

therefore for option b dencity is maximum......

ankit singh
3 years ago
 
 

DENSITY(d)= pM/RT,  i.e, it is directly proportional yo pressure and inversely proportional to temperatutre. Thus the condition is highest pressure and low temperature

like my answer

 
 

PV =nRT            ( n is number of moles of gas)

n = mass of gas /M.wt   & dencity(d) is mass/V

PV=(mass) RT/M.wt

P=dRT/M.wt         or

d= p(M.wt)/RT

from above eq , if the pressure is more and temprature is less then dencity will be high.....

therefore for option b dencity is maximum